# Evaluating the Evolution of YOLO (You Only Look Once) Models: A Comprehensive Benchmark Study of YOLO11 and Its Predecessors

### v1

@@ Abstract
This study presents a comprehensive benchmark analysis of various YOLO (You Only Look Once) algorithms, from YOLOv3 to the newest addition. It represents the first research to comprehensively evaluate the performance of YOLO11, the latest addition to the YOLO family. It evaluates their performance on three diverse datasets: Traffic Signs (with varying object sizes), African Wildlife (with diverse aspect ratios and at least one instance of the object per image), and Ships and Vessels (with small-sized objects of a single class), ensuring a comprehensive assessment across datasets with distinct challenges. To ensure a robust evaluation, we employ a comprehensive set of metrics, including Precision, Recall, Mean Average Precision (mAP), Processing Time, GFLOPs count, and Model Size. Our analysis highlights the distinctive strengths and limitations of each YOLO version. For example: YOLOv9 demonstrates substantial accuracy but struggles with detecting small objects and efficiency whereas YOLOv10 exhibits relatively lower accuracy due to architectural choices that affect its performance in overlapping object detection but excels in speed and efficiency. Additionally, the YOLO11 family consistently shows superior performance in terms of accuracy, speed, computational efficiency, and model size. YOLO11m achieved a remarkable balance of accuracy and efficiency, scoring mAP50-95 scores of 0.795, 0.81, and 0.325 on the Traffic Signs, African Wildlife, and Ships datasets, respectively, while maintaining an average inference time of 2.4ms, a model size of 38.8Mb, and around 67.6 GFLOPs on average. These results provide critical insights for both industry and academia, facilitating the selection of the most suitable YOLO algorithm for diverse applications and guiding future enhancements.
@@
YOLO (You Only Look Once), YOLO11, YOLOv10, Object detection, Ultralytics, Benchmark Analysis.I Introduction and Positioning
Object detection is an essential component of computer vision systems, enabling automated identification and localization of objects within images or video frames [34]. Its applications span from autonomous driving and robotics [16] [5] [20] [56] to inventory management, video surveillance, and sports analysis [4] [23] [55] [69].
@@
Following YOLOv3, the model’s development branched into various communities, leading to several notable iterations. YOLOv4 [6], developed by Bochkovskiy et al., introduced enhancements such as Spatial Pyramid Pooling (SPP) and the Path Aggregation Network (PAN). SPP aggregates features from multiple scales, preserving spatial information, while PAN improves the fusion of features between layers, resulting in improved speed and accuracy. YOLOv5 [60] marked a significant transition by moving from the Darknet framework to PyTorch, a popular deep learning library. This transition made the model more accessible and easier to customize. The architecture incorporated strided convolution layers, which reduced computational load, and Spatial Pyramid Pooling Fast (SPPF) layers, optimizing memory usage while maintaining high performance. YOLOv6 and YOLOv7 continued this trajectory with innovative architectures. YOLOv6 [29] introduced RepVGG, an architecture that simplified convolutional layers during inference, and CSPStackRep blocks, which improve accuracy by splitting the feature map into two parts to process them separately. In addition, YOLOv6 employed a hybrid channel strategy for better feature representation. YOLOv7 [65] leveraged the Extended Efficient Layer Aggregation Network (E-ELAN), a novel architecture that improved efficiency and effectiveness by enhancing information flow between layers.
@@
The most recent versions of YOLO, including YOLOv8, YOLOv9, YOLOv10, and YOLO11 represent the forefront of the model’s development. YOLOv8 [58], released by Ultralytics, introduced semantic segmentation capabilities, allowing the model to classify each pixel of an image, and provided scalable versions to meet various application needs, from resource-constrained environments to high performance systems alongside other tasks such as pose estimation, image classification, and oriented object detection (OOB). YOLOv9 [66] built on its predecessors’ architectural advancements with Programmable Gradient Information (PGI), which optimizes gradient flow during training, and the Generalized Efficient Layer Aggregation Network (GELAN), which further improved performance by enhancing layer information flow. YOLOv10 [64], developed by Tsinghua University, eliminated the need for Non-Maximum Suppression (NMS) used by its predecessors, a technique used to eliminate duplicate predictions and pick the bounding boxes with the most confidence, by introducing a dual assignment strategy in its training protocol. Additionally, YOLOv10 features lightweight classification heads, spatial-channel decoupled downsampling, and rank-guided block design, making it one of the most efficient and effective YOLO models to date. Lastly, YOLO11 [26], also introduced by Ultralytics, retains the capabilities of YOLOv8 with applications such as Instance Segmentation, Pose Estimation, and Oriented Object Detection while providing 5 scalable versions for different use cases. YOLO11 replaces the C2f block from YOLOv8 with the more efficient C3k2 block, delivering improved performance without compromising speed. Additionally, it introduces the C2PSA (Cross Stage Partial with Spatial Attention) module, which improves spatial attention in feature maps, increasing accuracy, especially for small and overlapping objects.
@@
This object detection algorithm has undergone several developments as seen in Figure 1 achieving competitive results in terms of accuracy and speed, making it the preferred algorithm in various fields such as ADAS (Advanced Driver-Assist System) [47], video surveillance [38], face detection [39], and many more [18]. For instance, YOLO plays a crucial role in the agriculture field by being implemented in numerous applications such as crop classification [1] [17], pest detection [33], automated farming [67] [37], and virtual fencing [62]. Moreover, YOLO has been utilized on numerous occasions in the field of healthcare such as cancer detection [cancer] [45], ulcer detection [2], medicine classification [36] [42], and health protocols enforcement [11].
@@
In recent years, Ultralytics has played a crucial role in the advancement of YOLO by maintaining, improving, and making these models more accessible [46]. Notably, Ultralytics has streamlined the process of fine-tuning and customizing YOLO models, a task that was considerably more complex in earlier iterations. The introduction of user-friendly interfaces, comprehensive documentation, and pre-built modules has greatly simplified essential tasks such as data augmentation, model training, and evaluation. Moreover, the development of scalable model versions allows users to select models tailored to specific resource constraints and application requirements, thereby facilitating more effective fine-tuning. For instance, YOLOv8n is favorable over YOLOv8m in scenarios where speed and computational efficiency are prioritized over accuracy, making it ideal for resource-constrained environments. The integration of advanced tools for hyperparameter tuning, automated learning rate scheduling, and model pruning has further refined the customization process. Continuous updates and robust community support have also contributed to making YOLO models more accessible and adaptable for a wide range of applications.
@@
This paper aims to present a comprehensive comparative analysis of the YOLO algorithm’s evolution. It makes a significant contribution to the field by offering the first comprehensive evaluation of YOLO11, the newest member of the YOLO family. By leveraging pre-trained models and fine-tuning them, we evaluate their performance across three diverse custom datasets, each with varying sizes and objectives. Consistent hyperparameters are applied to ensure a fair and unbiased comparison. The analysis delves into critical performance metrics, including speed, efficiency, accuracy, and computational complexity, as measured by GFLOPs count and model size. In addition, we explore the real-world applications of each YOLO version, highlighting their strengths and limitations across different use cases. Through this comparative study, we aim to provide valuable insights for researchers and practitioners, offering a deeper understanding of how these models can be effectively applied in various scenarios.
@@
The rest of this paper is organized as follows: Section 2 covers related work. Section 3 describes the datasets, the models, and the experimental setup, including the hyperparameters and evaluation metrics used. Section 4 presents the experimental results and comparative analysis alongside a discussion. Finally, Section 5 concludes with insights drawn from the study.
@@ II Related Work
The paper in [14] examines seven semantic segmentation and detection algorithms, including YOLOv8, for cloud segmentation from remote sensing imagery. It conducts a benchmark analysis to evaluate their architectural approaches and identify the most performing ones based on accuracy, speed, and potential applications. The research aims to produce machine learning algorithms that can perform cloud segmentation using only a few spectral bands, including RGB and RGBN-IR combinations.
@@
The authors of the paper in [22] review the evolution of the YOLO variants from version 1 to version 8, examining their internal architecture, key innovations, and benchmarked performance metrics. However, YOLOv9, YOLOv10, and YOLO11 are not considered in the analysis. The paper highlights the models’ applications across domains like autonomous driving and healthcare and proposes incorporating federated learning to improve privacy, adaptability, and generalization in collaborative training. The review, however, limits its focus to mAP (mean Average Precision) for accuracy evaluation, neglecting other key metrics such as Recall and Precision. Additionally, it considers FPS (frames per second) as the sole measure of computational efficiency, excluding the impact of preprocessing, inference, postprocessing times, GFLOPs, and size.
@@
The paper in [12] thoroughly analyzes single-stage object detectors, particularly YOLOs from YOLOv1 to YOLOv4, with updates to their architecture, performance metrics, and regression formulation. Additionally, it provides an overview of the comparison between two-stage and single-stage object detectors, several YOLO versions from version 1 to version 4, applications utilizing two-stage detectors, and future research prospects.
@@
The authors of the paper in [53] explore the evolution of the YOLO algorithms from version 1 to 10, highlighting their impact on automotive safety, healthcare, industrial manufacturing, surveillance, and agriculture. The paper highlights incremental technological advances and challenges in each version, indicating a potential integration with multimodal, context-aware, and General Artificial Intelligence systems for future AI-driven applications. However, the paper does not include a benchmarking study or a comparative analysis of the YOLO models, leaving out performance comparisons across the versions.
@@
The paper in [61] explores the development of the YOLO algorithm till the fourth version. It highlights its challenges and suggests new approaches, highlighting its impact on object detection and the need for ongoing study.
@@
This paper makes several key contributions: (i) It pioneers a comprehensive comparison of YOLO11 against its predecessors across their scaled variants from nano- to extra-large; (ii) it offers deep insights into the structural evolution of these algorithms by evaluating their performance across three diverse datasets of various object properties; and (iii) our performance evaluation extends beyond mAP and FPS to include critical metrics such as Precision, Recall, Preprocessing, Inference, and Postprocessing Time, GFLOPs, and model size. These metrics provide valuable insights to guide the selection of the optimal YOLO algorithm for specific use cases for both industry professionals and academics.
@@ III-A Datasets
This paper is made possible using several publicly accessible datasets on Kaggle and Roboflow. The selection of the datasets is based on the increasing implementation of the YOLO algorithms in the fields of Autonomous driving [47] [54] [30] [7], satellite imagery [31] [8] [44], and wildlife conservation [50] [68] [51]. Moreover, each picked dataset presents unique difficulties and situations for object detection with varying image sizes and number of observations alongside the number of classes.
@@ III-A1 Traffic Signs Dataset
The Traffic Signs dataset by Radu Oprea is an open-source dataset on Kaggle that contains around 55 classes across 3253 training and 1128 validation images of traffic signs in different sizes and environments [40]. All of the images in the dataset are initially in a size of 640640 with no labels for False Positives detection. Undersampling techniques were applied to this dataset to balance the different classes. This dataset is crucial for applications in autonomous driving, traffic management, road safety, and intelligent transportation systems. However, it presents challenges due to the varying sizes of target objects and the similarities in patterns across different classes, complicating the detection process.
@@ III-A2 Africa Wild Life Dataset
The Africa Wild Life dataset is an open-source Kaggle dataset by Bianca Ferreira, designed for real-time animal detection in nature reserves [59]. It features four common African animal classes: Buffalo, elephant, rhino, and zebra. Each class is represented by at least 376 images collected via Google image searches and manually labeled in the YOLO format. The challenges of this dataset are the varying aspect ratios, with each image containing at least one instance of the specified animal class and potentially multiple instances or occurrences of other classes. Moreover, overlapping these target objects makes the detection process more challenging. This dataset is essential for applications in wildlife conservation, anti-poaching efforts, biodiversity monitoring, and ecological research.

### v2

# YOLO Evolution: A Comprehensive Benchmark and Architectural Review of YOLOv12, YOLO11, and Their Previous Versions

@@ Abstract
This study presents a comprehensive benchmark analysis of various YOLO (You Only Look Once) algorithms. It represents the first comprehensive experimental evaluation of YOLOv3 to the latest version, YOLOv12, on various object detection challenges. The challenges considered include varying object sizes, diverse aspect ratios, and small-sized objects of a single class, ensuring a comprehensive assessment across datasets with distinct challenges. To ensure a robust evaluation, we employ a comprehensive set of metrics, including Precision, Recall, Mean Average Precision (mAP), Processing Time, GFLOPs count, and Model Size. Our analysis highlights the distinctive strengths and limitations of each YOLO version. For example: YOLOv9 demonstrates substantial accuracy but struggles with detecting small objects and efficiency whereas YOLOv10 exhibits relatively lower accuracy due to architectural choices that affect its performance in overlapping object detection but excels in speed and efficiency. Additionally, the YOLO11 family consistently shows superior performance maintaining a remarkable balance of accuracy and efficiency. However, YOLOv12 delivered underwhelming results, with its complex architecture introducing computational overhead without significant performance gains. These results provide critical insights for both industry and academia, facilitating the selection of the most suitable YOLO algorithm for diverse applications and guiding future enhancements.
@@
YOLO (You Only Look Once), YOLOv12, YOLO11, Object detection, Ultralytics, Benchmark Analysis.I Introduction
Object detection is an essential component of computer vision systems, enabling automated identification and localization of objects within images or video frames [37]. Its applications span from autonomous driving and robotics [17, 5, 21] [59] to inventory management, video surveillance, and sports analysis [4, 24, 58, 74].
@@
Following YOLOv3, the model’s development branched into various communities, leading to several notable iterations. YOLOv4 [7], developed by Bochkovskiy et al., introduced enhancements such as Spatial Pyramid Pooling (SPP) and the Path Aggregation Network (PAN). SPP aggregates features from multiple scales, preserving spatial information, while PAN improves the fusion of features between layers, resulting in improved speed and accuracy. YOLOv5 [64] marked a significant transition by moving from the Darknet framework to PyTorch, a popular deep learning library. This transition made the model more accessible and easier to customize. The architecture incorporated strided convolution layers, which reduced computational load, and Spatial Pyramid Pooling Fast (SPPF) layers, optimizing memory usage while maintaining high performance. YOLOv6 and YOLOv7 continued this trajectory with innovative architectures. YOLOv6 [31] introduced RepVGG, an architecture that simplified convolutional layers during inference, and CSPStackRep blocks, which improve accuracy by splitting the feature map into two parts to process them separately. In addition, YOLOv6 employed a hybrid channel strategy for better feature representation. YOLOv7 [69] introduced the Extended Efficient Layer Aggregation Network (E-ELAN) to improve efficiency and effectiveness by enhancing information flow between layers.
@@
The most recent versions of YOLO, including YOLOv8, YOLOv9, YOLOv10, YOLO11, and YOLOv12 represent the forefront of the model’s development. YOLOv8 [61], released by Ultralytics, introduced semantic segmentation capabilities, allowing the model to classify each pixel of an image, and provided scalable versions to meet various application needs, from resource-constrained environments to high performance systems alongside other tasks such as pose estimation, image classification, and oriented object detection (OOB). YOLOv9 [70] built on its predecessors’ architectural advancements with Programmable Gradient Information (PGI), which optimizes gradient flow during training, and the Generalized Efficient Layer Aggregation Network (GELAN), which further improved performance by enhancing layer information flow. YOLOv10 [68], developed by Tsinghua University, eliminated the need for Non-Maximum Suppression (NMS) used by its predecessors, a technique used to eliminate duplicate predictions and pick the bounding boxes with the most confidence, by introducing a dual assignment strategy in its training protocol. Additionally, YOLOv10 features lightweight classification heads, spatial-channel decoupled downsampling, and rank-guided block design, making it one of the most efficient and effective YOLO models to date. YOLO11 [28], also introduced by Ultralytics, retains the capabilities of YOLOv8 with applications such as Instance Segmentation, Pose Estimation, and Oriented Object Detection while providing 5 scalable versions for different use cases. YOLO11 replaces the C2f block from YOLOv8 with the more efficient C3k2 block, delivering improved performance without compromising speed. Additionally, it introduces the C2PSA (Cross Stage Partial with Spatial Attention) module, which improves spatial attention in feature maps, increasing accuracy, especially for small and overlapping objects. Lastly, YOLOv12 joins the YOLO family by also implementing applications such as Pose Estimation, Instance Segmentation, and Oriented Object Detection. Additionally, it provides 5 scalable version and introduces new architectural advancements such as the Area Attention (A2) and the Residual Efficient Layer Aggregation Networks (R-ELAN), which improve training stability and model convergence through block-level residual design and optimized feature aggregation.
@@
This object detection algorithm has undergone several developments, as seen in Figure reffig:yoloevolution achieving competitive results in terms of accuracy and speed, making it the preferred algorithm in various fields such as ADAS (Advanced Driver-Assist System) [50], video surveillance [41], face detection [42], and many more [19]. For instance, YOLO plays a crucial role in the agriculture field by being implemented in numerous applications such as crop classification [1] [18], pest detection [36], automated farming [71] [40], and virtual fencing [66]. Moreover, YOLO has been utilized on numerous occasions in the field of healthcare such as cancer detection [48], ulcer detection [2], medicine classification [39] [45], and health protocols enforcement [12].
@@
In recent years, Ultralytics has played a crucial role in the advancement of YOLO by maintaining, improving, and making these models more accessible [49]. Notably, Ultralytics has streamlined the process of fine-tuning and customizing YOLO models, a task that was considerably more complex in earlier iterations. The introduction of user-friendly interfaces, comprehensive documentation, and pre-built modules has greatly simplified essential tasks such as data augmentation, model training, and evaluation. Moreover, the development of scalable model versions allows users to select models tailored to specific resource constraints and application requirements, thereby facilitating more effective fine-tuning. For instance, YOLOv8n is favorable over YOLOv8m in scenarios where speed and computational efficiency are prioritized over accuracy, making it ideal for resource-constrained environments. The integration of advanced tools for hyperparameter tuning, automated learning rate scheduling, and model pruning has further refined the customization process. Continuous updates and robust community support helped YOLO models to be more accessible and adaptable for a wide range of applications.
@@
This paper aims to present a comprehensive comparative analysis of the YOLO algorithm’s evolution. It makes a significant contribution to the field by offering the first comprehensive evaluation of YOLOv12, the newest member of the YOLO family. By leveraging pre-trained models and fine-tuning them, we evaluate their performance across three diverse custom datasets, each with varying sizes and objectives. Consistent hyperparameters are applied to ensure a fair and unbiased comparison. The analysis focuses on numerous performance metrics, including speed, efficiency, accuracy, and computational complexity, as measured by GFLOPs count and model size. In addition, we explore the real-world applications of each YOLO version, highlighting their strengths and limitations across different use cases. Through this comparative study, we aim to provide valuable insights for researchers and practitioners, offering a deeper understanding of how these models can be effectively applied in various scenarios.
@@
The rest of this paper is organized as follows: Section 2 covers related work. Section 3 describes the datasets, the models, and the experimental setup, including the hyperparameters and evaluation metrics used. Section 4 discusses the experimental results and comparative analysis. Finally, Section 5 concludes with insights drawn from the study.
@@ II Related Work
The paper [15] examines seven semantic segmentation and detection algorithms, including YOLOv8, for cloud segmentation from remote sensing imagery. It conducts a benchmark analysis to evaluate their architectural approaches and identify the most performing ones based on accuracy, speed, and potential applications. The research aims to produce machine learning algorithms with cloud segmentation using a few spectral bands, including RGB and RGBN-IR combinations.
Showcasing the implementation of YOLO in marine ship detection, the study [25] propose the Adaptive Multi-scale YOLO (AM YOLO) to tackle marine ship segmentation challenges, introducing modules for enhanced detail extraction and cross-scale integration. While achieving up to 4.0% precision gains over YOLOv5s, the study focuses primarily on marine environments, limiting generalization to other domains—a gap addressed in this paper through broader dataset evaluations and versatile benchmarks alongside implementing newer state-of-the-art YOLO models.
In the paper [72], the authors developed a lightweight YOLO model optimized from YOLOv5, designed for multi-target garbage detection in dense and resource-constrained scenarios including MobileViTv3 for feature extraction, optimized MobileViT blocks for enhanced discrimination, and feature reuse techniques to minimize information loss. However, their reliance on YOLOv5 as a base architecture, which is less efficient and accurate than newer versions like YOLOv10 or YOLO11, limits both computational efficiency and performance—gaps addressed in this paper by leveraging state-of-the-art YOLO models and broader domain evaluations.
In the field of Smart Cities, the authors [6] work on a comprehensive survey of 3D object detection methods in road scene understanding. Their systematization offers a practical framework for comparing approaches, aiding researchers and practitioners in navigating the field. However, the study primarily focuses on other object detection models neglecting the use of YOLO in 3D object detection.
The paper [35] introduce the Efficient Attention Pyramid Transformer (EAPT), which enhances vision tasks by using Deformable Attention and the Encode-Decode Communication module (En-DeC) to capture more dynamic, global attention. While EAPT improves performance in various vision tasks, its integration with YOLO models is not explored. Incorporating these techniques into YOLO could enhance its feature extraction and global context understanding further improving the performance of YOLO variants. However, the study neglects the implementation of YOLO as a backbone, overlooking its potential in this area.
@@
The authors of [23] review the evolution of the YOLO variants from version 1 to version 8, examining their internal architecture, key innovations, and benchmarked performance metrics. The paper highlights the models’ applications across domains like autonomous driving and healthcare, and proposes incorporating federated learning to improve privacy, adaptability, and generalization in collaborative training. The review, however, limits its focus to mAP (mean Average Precision) for accuracy evaluation, neglecting other key metrics such as Recall and Precision. Additionally, it considers FPS (frames per second) as the sole measure of computational efficiency, excluding the impact of preprocessing, inference, postprocessing times, GFLOPs, and size.
@@
The paper [13] thoroughly analyzes single-stage object detectors, particularly YOLOs from YOLOv1 to YOLOv4, with updates to their architecture, performance metrics, and regression formulation. Additionally, it provides an overview of the comparison between two-stage and single-stage object detectors and applications utilizing two-stage detectors. However, not including newer YOLO models limits its comprehensiveness, leaving a gap in understanding the advancements and improvements introduced in more recent versions.
@@
The authors of [56] explore the evolution of the YOLO algorithms from version 1 to 10, highlighting their impact on automotive safety, healthcare, industrial manufacturing, surveillance, and agriculture. The paper highlights incremental technological advances and challenges in each version, indicating a potential integration with multimodal, context-aware, and General Artificial Intelligence systems for future AI-driven applications. However, the paper does not include a benchmarking study or a comparative analysis of the YOLO models, leaving out performance comparisons across the versions.
@@
The paper in [65] explores the development of the YOLO algorithm till the fourth version. It highlights its challenges and suggests new approaches, highlighting its impact on object detection and the need for ongoing study. However, it does not include newer YOLO models and lacks a comparative benchmark, limiting its ability to assess the progress and performance differences across various versions.
@@
This paper makes several key contributions: (i) It pioneers a comprehensive comparison of YOLOv12 against its predecessors across their scaled variants from nano- to extra-large; (ii) It provides deep insights into the architectural advancements of YOLO by analyzing key structural developments across versions. (iii) It evaluates YOLO models using three diverse datasets, reflecting various object properties and real-world applications, including Smart Cities, Satellite Imaging, and Wildlife Conservation. (iv) The performance evaluation extends beyond mAP and FPS, incorporating critical metrics such as Precision, Recall, Preprocessing Time, Inference Time, Postprocessing Time, GFLOPs, and model size. (v) These comprehensive metrics provide valuable insights for selecting the most optimal YOLO models, benefiting both industry professionals and academics. (vi) It offers specific use case recommendations, identifying the most suitable YOLO models for different scenarios and environments, such as resource-constrained deployments, real-time applications, and the detection of small or overlapping objects.
@@ III-A Datasets
This paper is made possible using several publicly accessible datasets on Kaggle and Roboflow. The selection of the datasets is based on the increasing implementation of the YOLO algorithms in the fields of Autonomous driving [50] [57] [32] [8], satellite imagery [33] [9] [47], and wildlife conservation [53] [73] [54]. Moreover, each picked dataset presents unique difficulties and situations for object detection with varying image sizes and number of observations alongside the number of classes to ensure a comprehensive evaluation.
@@ III-A1 Traffic Signs Dataset
The Traffic Signs dataset by Radu Oprea is an open-source dataset on Kaggle containing around 55 classes across 3253 training and 1128 validation images of traffic signs in various sizes and environments [43]. All images in the dataset are initially sized 640640, with no labels for False Positives detection. To balance the classes, undersampling techniques were applied, as shown in Figure 2. After preprocessing the dataset by removing classes with low observation counts, 24 classes remained, with 3233 images split into 70% training, 20% validation, and 10% testing, with no data augmentation techniques applied. This dataset is vital for applications in autonomous driving, traffic management, road safety, and intelligent transportation systems. Additionally, it presents challenges due to the varying sizes of target objects and the similarities in patterns across different classes, which complicate the detection process.
@@ III-A2 Africa Wild Life Dataset
The Africa Wildlife dataset is an open-source Kaggle dataset by Bianca Ferreira, designed for real-time animal detection in nature reserves [62]. It features four common African animal classes: Buffalo, elephant, rhino, and zebra. Each class is represented by at least 376 images, which were collected via Google image searches without any data augmentation techniques applied. The dataset is split into 70% training, 20% validation, and 10% testing, with all images manually labeled in the YOLO format, as shown in Figure 3. The dataset presents challenges such as varying aspect ratios, with each image containing at least one instance of the specified animal class and potentially multiple occurrences of other classes. Furthermore, the overlapping of target objects increases the difficulty of detection. This dataset is crucial for applications in wildlife conservation, anti-poaching efforts, biodiversity monitoring, and ecological research.